

********************************************************************
UARS WEEKLY STATUS REPORT - 16 FEB 2001
********************************************************************
***********
CDHF
***********
*************IMPORTANT*************
UARS CDHF REHOST:
_ Level 0 and Quicklook are being received on both the rehost and current
system. This will continue until NRT implementation is resolved.
- To accomodate concerns for flight safety, we have moved the NRT
software distribution back to CDHF2 .
NRT will stay on CDHF2 until confirmation from the instrumentors
has been received. PI teams need to migrated their
NRT software that is located at GSFC current system or on the RAC, to use
data flowing from the rehost machine (CDHFA/CDHFB). Please contact PAC for
any help you may need.
The lack of true-redundancy on the Level-0 processor inhibits the ability
to send NRT to both computers at the same time, thus the decision.
- PEM V3.32 (JOBS 1 and 2) have both tested successfully on the
Rehost System. Waiting and working with PEM for validation results
_________________
o DAAC DATA:
- Data created from 18-Jan-2001 through 7-Feb-2001 for MLS V5 has been
loaded into the DAAC Transfer Tables. DAAC has been notified.
- Data received and cataloged in the month of December 2000 has been
previewed in preparation for sending to the DAAC.
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
ACRIM II continues to function normally and gather high quality data.
Comparisons with ACRIM III demonstrate the close agreement of both
sets of observations relative to solar variable features associated with
solar maxima
Re-processing using ACRIMSAT/ACRIM III algorithms is proceeding.
The goal is to release re-processed ACRIM II results by 4/1/01. These
version 2 results will be posted at the LaRC and GSFC DAAC's and on
the official ACRIM web page: ACRIM.COM.
***********
PEM
***********
PEM/AXIS was activated in its full power mode on January 11. It was
change to half power mode on January 28 when the AXIS-2 unit was deactivated.
AXIS-1 was again active in the half power mode between January 28 and
February 7. At the present time, energy deposition from PEM/AXIS is not being
computed due to a software problem which PEM is to repair.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. Initial tests of telemetry transfers through
PEM2K:: have been successful; however, the source of these failures has been
traced to quota limits and these have been corrected. PEM Telemetry is runing
in duplicate streams through PEM:: and PEM2K::. We find that files which
transfer from the CDHF to PEM:: and PEM2K:: are now identical. PEM has
requested DAP command capability with the CMS from PEM2K::.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to function normally, making wind measurements during
the sunlit parts of orbits nearly every day, and measuring nighttime winds
when power is available.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in an Incident Review Board meeting on
16 February which was held as part of the anomaly investigation>
process in regard to the SOLSTICE load uplinking incident on
13 December 2000.
Daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is now complete
to 15 March 2001.
***********
FOT
***********
REPORTING PERIOD
This report covers 03 February 2001 (Orbit 51376, GMT Day 034) through 09
February 2001 (Orbit 51480, GMT Day 040).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 028/1250 51293 Power On
A Reverse-to-Forward Yaw Around maneuver (number 96) was successfully executed
in Orbit 51477 beginning 040/17:42:21 and completing at 040/18:19:31 on 09
February 2001. The SOLSTICE instrument was in day only and all other
instruments were in their off or standby mode (including HRDI and PEM). Due to
the SA position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 54.1 minutes. Post maneuver analysis of
data indicate that during the Yaw Around, the end-of-night load bus voltage was
24.64 volts and the average depths of discharge (DODs) on Batteries 2 and 3
were approximately 21.5 and 22.0 percent respectively. The next (forward-to-
reverse) YAW is scheduled for 15 March 2001 (DOY 074).
The beta angle ranged between the angles of 28.6 degrees to 3.9 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 48 degrees on 26 February 2001 (DOY 057).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.9 and 1.1 degrees C. Additional battery performance data for each
battery on 02 February 2001 and 09 February 2001 are provided in Appendix D.
The clock error ranged from Ð13.9 to 10.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 21
¥ HALOE = 01
¥ SUSIM = 01
¥ ACRIM = 01
¥ PEM = 02
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 04
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/027 - 2287.499274
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculation supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/039 show the observatory data loss to be 300 hours, 13
minutes, 42.088 seconds. This shows an overall increase of 15.040 seconds since
the last report period This is a 0.39 percent data loss, which equals a 99.61
percent data capture for the mission.
The increase of 15.040 seconds occurred over one (1) day:
· 15.040 seconds on Day 039 due to a Data Hit
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
After completing the 96th yaw maneuver on UARS successfully, the first FSS
update showed approximately 300 arc sec residuals. The FOT was pleased with
this result and continues to make adjustments to the TAM alignment to improve
its performance and UARS's attitude.
System administrators replaced the bad hard disk on the DFLX system. No
further problems have been reported regarding the CMS to DFLX file transfer or
with the DFLX load blocking process.
MEETINGS
The FOT attended a Solstice Meeting on 06 February 2001.
The FOT attended a Flight Software Design Review on 08 February 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 15 March 2001 (DOY 074). This maneuver
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, O
tions
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 09 February 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
51376 034/0137 01-074 OBC/CLK Clk Adj 0.00 to 20.6 msec/day
51378 034/0501 01-073 NBTR NBTR PBK/RTS-58
51379 034/0641 01-073 NBTR NOOPS D-034
51389 034/2215 01-075 NBTR NOOPS D-035
51394 035/0602 01-075 NBTR NBTR PBK/RTS-58
51394 035/0602 01-076 OBC/CLK Clk Adj 20.6 to 0.00 msec/day
51395 035/0642 01-075 NBTR NOOPS D-035
51404 035/2140 01-077 NBTR NOOPS D-036
51404 035/2140 01-078 OBC/CLK Clk Adj 0.00 to -20.6 msec/day
51408 036/0605 01-077 NBTR NBTR PBK/RTS-58
51409 036/0644 01-077 OBC NOOPS D-036
51415 036/1524 01-079 ACS TAM CAL
51420 036/2322 01-081 HALOE Power OFF
51420 037/0004 01-080 NBTR NBTR-B STOP REC.
51421 037/0141 01-082 OBC/CLK Clk Adj -20.6 to 0.00 msec/day
51422 037/0239 01-080 NBTR NBTR PBK
51431 037/1625 01-083 MACS TAM CAL
51445 038/1444 01-085 OBC OBC PATCH FOR ESA
51450 038/2324 01-084 PEM AXIS OFF
51460 039/1526 01-086 ACS TAM Calibration
51465 039/2324 01-087 SUSIM Power OFF
51465 039/2324 01-088 ACRIM Power OFF
51465 039/2324 01-089 PEM Power OFF
51476 040/1708 01-090 OBC Gyro Bias (yaw) FSS/TAM Mon
51476 040/1708 01-091 OBC Yaw Maneuver
51477 040/1845 01-092 OBC FSS/TAM to update mode
___________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 09 February 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 09 February 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
51376 034/0152 Clk Adj 0.00 to 20.6 msec/day
51394 035/0607 Clk Adj 20.6 to 0.00 msec/day
51404 035/2140 Clk Adj 0.00 to -20.6 msec/day
51421 037/0144 Clk Adj -20.6 to 0.00 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 09 February 2001
02 Feb 2001 - GMT Day 033 Beta = 32.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ---------- ---------- ------ ------- ------
1 +039.2/+000.0 -3.68/-4.62 +0.0/+0.0 13.0 0.000 0.0
2 +11.2/-22.4 +5.47/+3.87 +28.8/-14.4 25.9 1.040 20.3
3 +22.4/-22.4 +4.51/+2.92 +28.4/-15.2 25.9 1.030 19.9
02 Feb 2001 - GMT Day 033 Beta = 32.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ----------- ------ ------- ------
1 +000.0/-028.0 -3.68/-4.94 0.0/+0.0 13.0 0.000 0.0
2 +16.8/-61.6 +5.15/+3.87 +29.6/-13.6 24.8 1.020 21.0
3 +33.6/-50.4 +4.19/+2.92 +28.8/-15.6 24.8 1.010 20.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
-no changes V/T 5
___________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 February 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-010 FOT N/A 034-035 N/A DFLX crashed; no transfer
01-009 FOT 51408 036/0606 14170 Generic Late Acq.
01-009 FOT 51472 040/1125 14170 Generic Late Acq.
01-009 FOT 51476 040/1632 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-005 FOT 51276 027/0917 N/A Load u027z101 cont. no commands
01-006 FOT 51280 027/1514 N/A SSPP missed one star track
01-007 FOT 51310 029/1614 N/A CMS no response to xfer
01-010 FOT N/A 034-035 N/A DFLX crashed; no xfer
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-009 FOT see list Feb 01 14170 Generic Late Acquisition
SEVNTFW01/06